OpenStudy (anonymous):

for a meddle first one to answer gets the meddle what is 0*0

OpenStudy (amoodarya):

0*(every number )=0

OpenStudy (amoodarya):

1234*0=0 or 764768374*0=0

OpenStudy (anonymous):

i know im not dum im just doing random questions

OpenStudy (amoodarya):

example make it easy to understand , lol don't care about what people say believing your self and go you own way